OpenStudy (anonymous):

I need help, because I don't know how to do this. I will medal!! Linda is planning to build a rectangular deck that is 12 feet longer than it is wide. Find its width if its area is 589 square feet. Write an equation that models this problem?

OpenStudy (anonymous):

length=l width=w l=w+12 l*w=area (w+12)*w=589 => w^2 +12w=589 =>w^2 +12w - 589=0

OpenStudy (anonymous):

\[w^{2} + 12w -589=0\]

OpenStudy (anonymous):

@annwesha Can you help me with a few more, if you are up to it? Thanks

OpenStudy (anonymous):

Yes, sure.. Anytime...